OpenStudy (anonymous):

Solve the equation... x^2(12x+101)/63=x

OpenStudy (anonymous):

clearly x=0 is a solution

OpenStudy (anonymous):

I was thinking that was an option...I just didn't know where to go from there:)

OpenStudy (lgbasallote):

you cant really assume that yet @furan143 ...youll have to transform and solve it first...

OpenStudy (anonymous):

Where do I start?

OpenStudy (anonymous):

after that you should divide the whole equation with x which leaves you with a quadratic equation which can be solved

OpenStudy (anonymous):

so x=9 or x=-7/12?

OpenStudy (anonymous):

I am still not seeing where you where able to break it all down to 9 and -7/12

OpenStudy (anonymous):

\[x^{2}(12x+101)/63=x\] so \[x^{2} (12x +101) = 63x\] so\[12x^{3}+101x^{2}=63x\] move 63x to the left of the equation \[12x^{3} +101x^{2} - 63 x = 0\]

OpenStudy (anonymous):

\[x(12x^{2} + 101 x - 63)=0\] so x(x+9)(12x-7)=0

OpenStudy (anonymous):

thus x=0 or x+9 = 0, or 12x-7 = 0 Then, x = 0 or x = -9 or x = 7/12